### v2

# Stiefel-Whitney Classes for Finite Symplectic Groups

Abstract.
Let $q$ be an odd prime power, and $G=\operatorname{Sp}(2n,q)$ the finite symplectic group. We give an expression for the total Stiefel-Whitney Classes (SWCs) for orthogonal representations $\pi$ of $G$, in terms of character values of $\pi$ at elements of order $2$. We give “universal formulas” for the fourth and eighth SWCs. For $n=2$, we compute the subring of the mod $2$ cohomology generated by the SWCs $w_{k}(\pi)$.
Key words and phrases:
Stiefel-Whitney classes, Symplectic Groups, Finite groups of Lie type, Weil Representations
1991 Mathematics Subject Classification:
Primary 20G40, Secondary 55R40
1. Introduction
Stiefel-Whitney Classes (SWCs) are interesting natural cohomological invariants of orthogonal representations. In this paper we present a formula for SWCs for the finite symplectic groups in odd characteristic.
This paper is part of a project to understand SWCs for finite groups of Lie type. Let $q$ be an odd prime power throughout. A formula determining SWCs for $G=\operatorname{GL}(n,q)$ was discovered in [GJgln], for $G=\operatorname{SL}(2,q)$ in [NSSL2] and for $G=\operatorname{SL}(2n+1,q)$ in [MSSLn]. In this paper we find similar expressions when $G=\operatorname{Sp}(2n,q)$.
Write $\operatorname{H}^{*}(G)$ for the mod $2$ cohomology $\operatorname{H}^{*}(G,\mathbb{F}_{2})$, and $\operatorname{H}^{*}_{\operatorname{SW}}(G)$ for the subalgebra of $\operatorname{H}^{*}(G)$ generated by SWCs of orthogonal representations. The diagonal matrices in $G$ with eigenvalues $\pm 1$ form a subgroup we denote by $Z_{X}$. (It is the center of another subgroup $X$ which we will encounter later.) The mod $2$ cohomology of $Z_{X}$ is a polynomial algebra:
$\operatorname{H}^{*}(Z_{X})\cong\mathbb{F}_{2}[v_{1},\ldots,v_{n}],$
where each $v_{i}$ is the first SWC of a certain linear character of $Z_{X}$; in particular it has degree 1. Let $\mathfrak{e}_{i}=v_{i}^{4}$, and let $\mathcal{E}_{k}$ be the $k$th elementary symmetric polynomial in the $\mathfrak{e}_{i}$.
Theorem 1.1.
The restriction map $\operatorname{H}^{*}_{\operatorname{SW}}(G)\to\operatorname{H}^{*}(Z_{X})$ is injective, and its image is contained in $\mathbb{F}_{2}[\mathcal{E}_{1},\ldots,\mathcal{E}_{n}]$.
Since no information is lost by restriction, we will express SWCs in terms of the $\mathcal{E}_{k}$. For $0\leq i\leq n$, let $g_{i}\in Z_{X}$ with 1-1 having multiplicity $2i$ as an eigenvalue, and $1$ having multiplicity 2(ni)2(n-i). For example $g_{0}=\mathbbm{1}$, the identity matrix. Given a representation $\pi$, write $\chi_{\pi}$ for its character. Here is our “universal” formula for the 4th and 8th SWCs:
Theorem 1.2.
If $\pi$ is an orthogonal representation of $\operatorname{Sp}(2n,q)$, then
w4(π)=18(degπχπ(g1))1 for all n,
[…]
and
$w_{8}(\pi)=r_{1}\mathcal{E}_{2}+\left(\binom{r_{1}}{2}+\binom{r_{2}}{2}\right)\mathcal{E}_{1}^{2}\quad\text{ for $n\geq 2$}.\\$
Here
$r_{1}=\frac{1}{16}(\deg\pi-\chi_{\pi}(g_{2}))$
and
$r_{2}=\frac{1}{16}(\deg\pi-2\chi_{\pi}(g_{1})+\chi_{\pi}(g_{2})).$
A general formula for the total SWCs $w(\pi)$ is found in Theorem 3.6, although it takes some computation to extract individual SWCs $w_{k}(\pi)$ from this. For example:
Theorem 1.3.
The total SWC of an orthogonal representation $\pi$ of $\operatorname{Sp}(4,q)$ is
$w(\pi)=((1+\mathfrak{e}_{1})(1+\mathfrak{e}_{2}))^{r_{\pi}}(1+\mathfrak{e}_{1}+\mathfrak{e}_{2})^{s_{\pi}},$
where
$r_{\pi}$ $=\frac{1}{16}\Big(\chi_{\pi}(\mathbbm{1})-\chi_{\pi}(-\mathbbm{1})\Big)\quad\quad\text{ and }$
$s_{\pi}$ $=\frac{1}{16}\Big(\chi_{\pi}(\mathbbm{1})+\chi_{\pi}(-\mathbbm{1})-2\chi_{\pi}(g_{1})\Big).$
We use this theorem to compute:
Corollary 1.3.1.
The subalgebra
$\operatorname{H}^{*}_{\operatorname{SW}}(\operatorname{Sp}(4,q))\cong\mathbb{F}_{2}[\mathfrak{e}_{1}+\mathfrak{e}_{2},\mathfrak{e}_{1}\mathfrak{e}_{2}].$
One can in principle produce universal formulas for all $w_{k}$, akin to those of Theorem 1.2 by the following. Let $G_{n}=\operatorname{Sp}(2n,q)$.
Theorem 1.4.
Let $m\geq n$. The restriction map $\iota_{n}^{*}:\operatorname{H}^{i}(G_{m})\to\operatorname{H}^{i}(G_{n-1})$ is injective for i<4n1i<4n-1.
(Compare [NSn, Corollary 6.7] and [GJgln, Theorems 3, 8].)
A representation $\pi$ of a group $G$ has a total Chern class $c(\pi)\in\operatorname{H}^{*}(G,\mathbb{Z})$. Under the coefficient map to $\operatorname{H}^{*}(G,\mathbb{F}_{2})$, these map to $w(\pi\oplus\pi^{\vee})$. Hence this “mod $2$ Chern class of $\pi$” is computable for $G=\operatorname{Sp}(2n,q)$ by our formulas.
The paper is laid out as follows. Preliminaries are reviewed and developed in Section 2. Section 3 contains the heart of the paper; we prove Theorem 1.1, and give the product formula for the total SWC, as Theorem 3.6. Some simplifications come when the representation is irreducible, by a formula of Gow. We develop our formulas for $\operatorname{Sp}(4,q)$ and $\operatorname{Sp}(8,q)$ in Section 4, and prove Theorem 1.3 and Corollary 1.3.1. We also illustrate application of our formula by computing the mod $2$ Chern class of the Weil representations. Finally, in Section 5 we establish Theorem 1.4, and deduce Theorem 1.2 from this. Certain technical arguments belonging to vector bundle theory are sketched in the Appendix.
2. Notations and Preliminaries
As this paper is a continuation of [NSSL2], we use the same notations and conventions, which we now review.
2.1. Representations
Let $G$ be a finite group. All the representations $(\pi,V)$ considered in this paper (before the Appendix) are complex finite dimensional. Let $\operatorname{Irr}(G)$ be the set of isomorphism classes of irreducible representations of $G$. Write $(\pi^{\vee},V^{\vee})$ for the dual representation. If $H$ is a subgroup of $G$, write $\operatorname{res}^{G}_{H}\pi$ or $\pi|_{H}$ for the restriction of $\pi$ to $H$. A linear character $\chi$ of $G$ is a degree 1 representation. We call $\chi$ quadratic when $\chi^{2}=1$.
We say $\pi$ is orthogonal (resp., symplectic), provided there exists a non-degenerate $G$-invariant symmetric (resp., antisymmetric) bilinear form $B:V\times V\to\mathbb{C}$. When $\pi$ is self-dual and irreducible, it is either orthogonal or symplectic. In this case, the Frobenius-Schur Indicator $\varepsilon(\pi)$ is a sign defined as $1$ when $\pi$ is orthogonal, and 1-1 when $\pi$ is symplectic. Whereas it is 0, when $\pi$ is not self-dual.
One can symmetrize a general $(\pi,V)$ by defining $S(\pi):=\pi\oplus\pi^{\vee}$ on the vector space $V\oplus V^{\vee}$. Under the symmetric $G$-invariant bilinear map $B$ on $V\oplus V^{\vee}$ as $B((v,\alpha),(w,\beta))=\langle\alpha,w\rangle+\langle\beta,v\rangle$, the representation $S(\pi)$ is orthogonal. We call $S(\pi)$ the symmetrization of $\pi$. (In [shl, Definition 4.3 ] it is called the hyperbolic space on $V$.)
Every orthogonal representation $\Pi$ of $G$ can be decomposed as
(1) $\Pi\cong\bigoplus_{i}\pi_{i}\oplus\bigoplus_{j}S(\varphi_{j}),$
such that each $\pi_{i}$ is irreducible orthogonal and $\varphi_{j}$ are irreducible non-orthogonal representations of $G$.
A representation $\pi$ of $G$ is said to be an orthogonally irreducible representation (OIR), provided $\pi$ is orthogonal, and can not be decomposed into a direct sum of orthogonal representations. An irreducible representation $\pi$ is orthogonally irreducible if and only if $\pi$ is orthogonal. Moreover, for $\varphi$ irreducible and non-orthogonal, its symmetrization $S(\varphi)$ is an OIR.
2.2. Detection
As in the earlier work, we make use of detection. To recall this notion, let $i:H\hookrightarrow G$ be a subgroup. We say $H$ detects the mod $2$ cohomology of $G$, when the restriction map
$i^{*}:\operatorname{H}^{*}(G)\rightarrow\operatorname{H}^{*}(H)$
is injective, and we say $H$ detects SWCs of $G$ when the restriction of $i^{*}$ to $\operatorname{H}^{*}_{\operatorname{SW}}(G)$ is injective. Often the cohomology of $H$ admits an easy description, e.g., when it is polynomial. Then it is convenient to give our formulas there, since no information is lost. Let $N_{G}(H)$ be the normalizer of $H$ in $G$, which acts on $H$ by conjugation. This induces an action of $N_{G}(H)$ on $\operatorname{H}^{*}(H)$, and generally the image of $i^{*}$ is contained in the subalgebra $\operatorname{H}^{*}(H)^{N_{G}(H)}$ fixed under this action.
2.3. Characteristic Classes
Let $\pi$ be an orthogonal representation of degree $d$. Associated to $\pi$ are cohomological invariants
$w_{i}(\pi)\in\operatorname{H}^{i}(G)\quad;\quad i=0,1,2,\ldots,d$
known as the $i$th Stiefel-Whitney Class (SWC) of $\pi$. Their sum $w(\pi)=w_{0}(\pi)+w_{1}(\pi)+\ldots$ is called the total SWC of $\pi$. We refer the reader to [NSSL2, Section 2.3] for detailed description.
Also, associated to a complex representation $\pi$ of $G$ are cohomology classes $c_{i}(\pi)\in\operatorname{H}^{2i}(G,\mathbb{Z}),$ called Chern classes (CCs). Their sum
$c(\pi)=c_{0}(\pi)+c_{1}(\pi)+c_{2}(\pi)+\ldots\in\operatorname{H}^{*}(G,\mathbb{Z})$
is called the total Chern class of $\pi$. We have
(2) $w(S(\pi))=\kappa(c(\pi)),$
where $\kappa:\operatorname{H}^{*}(G,\mathbb{Z})\rightarrow\operatorname{H}^{*}(G,\mathbb{F}_{2})$ is the coefficient homomorphism of cohomology. (See [ganguly2022stiefel, Lemma 1], based on [milnor, Problem 14-B] for proof.) So we interpret $w(S(\pi))$ as the “mod $2$ Chern class” of $\pi$.
For $n,i\geq 0$, there are additive homomorphisms on cohomology, called Steenrod Squares,
$\operatorname{Sq}^{i}:\operatorname{H}^{n}(G)\rightarrow\operatorname{H}^{n+i}(G).$
These operations are functorial, meaning for a group homomorphism $\varphi:G_{1}\to G_{2}$, we have
$\varphi^{*}(\operatorname{Sq}^{i}y)=\operatorname{Sq}^{i}(\varphi^{*}(y))\text{ for all }y\in\operatorname{H}^{i}(G_{2}).$
They satisfy $\operatorname{Sq}^{i}(x)=x\cup x$ for $i=\deg(x)$, and $\operatorname{Sq}^{i}(x)=0$ for $i>\deg(x)$. There is the Cartan Formula: For $x,y\in\operatorname{H}^{*}(G)$,
(3) $\operatorname{Sq}^{n}(x\cup y)=\sum_{i+j=n}(\operatorname{Sq}^{i}x)\cup(\operatorname{Sq}^{j}y).$
The well-known Wu’s formula states:
Proposition 2.1 ([may], Chapter 23, Section 6).
Let $\pi$ be an orthogonal representation of $G$. The cohomology class $\operatorname{Sq}^{i}(w_{j}(\pi))$ can be expressed as a polynomial in $w_{1}(\pi),\ldots,w_{i+j}(\pi)$:
$\operatorname{Sq}^{i}(w_{j}(\pi))=\sum_{t=0}^{i}{j+t-i-1\choose t}w_{i-t}(\pi)w_{j+t}(\pi).$
Corollary 2.1.1.
Suppose $w_{1}(\pi)=w_{2}(\pi)=0$. Then if $w_{i}(\pi)\neq 0$, then $i$ is a multiple of $4$.
Proof.
This is clear. ∎
2.4. Symmetric Functions
The elementary symmetric functions $\mathcal{E}_{k}(\mathbf{x})$ over $\mathbb{F}_{2}$ in variables $x_{1},x_{2},\ldots,x_{n}$ are defined for $k\leq n$ as,
$\mathcal{E}_{k}(\mathbf{x}):=\sum_{1\leq i_{1}<\ldots<i_{k}\leq n}x_{i_{1}}x_{i_{2}}\ldots x_{i_{k}}\in\mathbb{F}_{2}[x_{1},\ldots,x_{n}].$
Whereas for $k>n$, these are defined to be 0. For example, $\mathcal{E}_{1}(x_{1},x_{2})=x_{1}+x_{2}$ and $\mathcal{E}_{2}(x_{1},x_{2})=x_{1}x_{2}$. Of course,
$\prod_{i=1}^{n}(1+x_{i})=1+\mathcal{E}_{1}(\mathbf{x})+\ldots+\mathcal{E}_{n}(\mathbf{x}).$
2.5. Cyclic Groups
Let $k$ be even, and $G=C_{k}$ the cyclic group of order $k$. Let $g$ be a generator of $G$. We write $g^{k/2}=-1$, the unique order $2$ element of $G$. We say a linear character $\chi:G\to\mathbb{C}$ is odd, when $\chi(-1)=-1$, and even, when $\chi(-1)=1$. Let ‘sgn’ denote the unique non-trivial quadratic character of $G$.
It is known [KT] that for $k\equiv 2\pmod{4}$, $\operatorname{H}^{*}(C_{k})=\mathbb{F}_{2}[v]$ where $v=w_{1}(\operatorname{sgn})$. In this case, (see [NSSL2, Lemma 2.5] for instance) for a representation $\pi$,
$w(\pi)=(1+v)^{b_{\pi}}$
with $b_{\pi}=\cfrac{1}{2}\left(\deg\pi-\chi_{\pi}(-1)\right)$.
Let $C_{2}^{n}$ be the $n$-fold product of $C_{2}$, with projection maps $\operatorname{pr}_{i}:C_{2}^{n}\to C_{2}$ for $i=1,\ldots,n$. By Künneth, we have
(4) $\operatorname{H}^{*}(C_{2}^{n})=\mathbb{F}_{2}[v_{1},\dots,v_{n}],$
where we put $v_{i}=w_{1}(\operatorname{sgn}\circ\operatorname{pr}_{i})$ for $1\leq i\leq n$.
2.6. OIRs of a Direct Product of Groups
Let $G_{1},\ldots,G_{n}$ be finite groups and $G=G_{1}\times\cdots\times G_{n}$ be their direct product. Given $G_{i}$-representations $(\pi_{i},V_{i})$, one can form their external tensor product $\pi=\pi_{1}\boxtimes\cdots\boxtimes\pi_{n}$ from the action of the product group $G$ on the tensor space $V_{1}\otimes\cdots\otimes V_{n}$ as,
$(g_{1},\ldots,g_{n})(v_{1}\otimes\cdots\otimes v_{n})=g_{1}v_{1}\otimes\cdots\otimes g_{n}v_{n}$
for $(g_{1},\ldots,g_{n})\in G$ and $v_{1}\otimes\cdots\otimes v_{n}\in V_{1}\otimes\cdots\otimes V_{n}.$ (See [fulton, Exercise 2.36, page 24].) If each $\pi_{i}$ is irreducible, then so is $\pi$. All irreducible representations of $G$ decompose in this manner. The Frobenius-Schur indicator of $\pi$ is given by
(5) $\begin{split}\varepsilon(\pi)&=\varepsilon(\pi_{1})\varepsilon(\pi_{2})\ldots\varepsilon(\pi_{n}).\end{split}$
Write $F(\pi)$ for the multiset $\{\pi_{1},\pi_{2},\ldots,\pi_{n}\}$. Then from (5), we can describe the OIRs of $G$ in terms of irreducible representations of $G_{i}$ as follows:
  1. (1)
    Irreducible orthogonal representations of the form $\pi=\pi_{1}\boxtimes\cdots\boxtimes\pi_{n},$ where $\pi_{i}\in\operatorname{Irr}(G_{i})$ are self-dual for each $i$ and an even number of representations in $F(\pi)$ are symplectic.
  2. (2)
    Symmetrization of irreducible non-orthogonal representations
    $\pi=S(\varphi)=S(\varphi_{1}\boxtimes\cdots\boxtimes\varphi_{n}),$ where $\varphi_{i}\in\operatorname{Irr}(G_{i})$ for $i=1,\ldots,n$ and exactly one of the following holds:
    1. (a)
      At least one of $\varphi_{i}$ is not self-dual.
    2. (b)
      Each $\varphi_{i}$ is self-dual and there is an odd number of symplectic representations in $F(\varphi)$.
2.7. Quaternion Group
Let $Q$ be the quaternion group of order 8. There are four linear characters of $Q$, which we denote by $1$, $\chi_{1},\chi_{2},\chi_{1}\otimes\chi_{2}$. Each one is quadratic. The group also possesses a unique irreducible representation $\rho$ of degree $2$. It is symplectic.
Let $Q^{n}$ be the $n$-fold product of $Q$. From Section 2.6 above, any irreducible representation $\pi$ of $Q^{n}$ has the form $\pi\cong\pi_{1}\boxtimes\cdots\boxtimes\pi_{n}$ where each $\pi_{i}\in\operatorname{Irr}(Q)$. Since every representation of $Q$ is self-dual, the same is true for $Q^{n}$.
Put $I_{\pi}=\{i:\pi_{i}\cong\rho\}$, and $r(\pi)=|I_{\pi}|$. From (5) we obtain:
Lemma 2.2.
An irreducible representation $\pi$ of $Q^{n}$ is orthogonal if and only if $r(\pi)$ is even.
Let $Z$ be the center of $Q$, which is $\{\pm 1\}$. For a linear character $\theta$ of $Z^{n}$, write $\theta,\pi]$ for the multiplicity of $\theta$ in $\pi|_{Z^{n}}$.
Lemma 2.3.
Let $\pi$ be an orthogonal representation of $Q^{n}$. Let $\theta$ be a non-trivial linear character of $Z^{n}$. Then $\theta,\pi]$ is a multiple of $4$.
Proof.
We may assume that $\pi$ is an OIR. Let $r=r(\pi)$.
If $r=0$, then $\pi|_{Z}$ is trivial, since $Z$ is in the kernel of all linear characters of $Q$. So take $r>0$. Define a linear character $\theta_{\pi}$ of $Z^{n}$ by $\theta_{\pi}=\boxtimes_{i}\theta_{i}$, with
[…]
Then $\pi|_{Z^{n}}\cong 2^{r}\theta_{\pi}$, since $\rho|_{Z}=\operatorname{sgn}\oplus\operatorname{sgn}$. If $\pi$ is irreducible orthogonal, then $r$ is even by Lemma 2.2 and therefore $4$ divides $\theta_{\pi},\pi]$, and other $\theta,\pi]=0$.
Otherwise $\pi=S(\phi)$ for an irreducible symplectic representation $\phi$. In this case,
$\theta_{\pi},S(\phi)]=2^{r}+2^{r}=2^{r+1},$
which is again divisible by $4$ for $r>0$. ∎
2.8. $S_{n}$-invariant Representations of Elementary Abelian $2$-groups
Let $E$ be an elementary abelian $2$-group of rank $n$. View $E$ as an $\mathbb{F}_{2}$-vector space, say with basis $e_{1},\ldots,e_{n}$. For $e\in E,$ put $|e|=\#\{i:c_{i}=1\}$ when $e=\sum\limits_{i=1}^{r}c_{i}e_{i}$ with $c_{i}\in\mathbb{F}_{2}$. Put $\mathcal{O}_{k}=\{e\in E:|e|=k\}.$
Let $E^{\vee}=\operatorname{Hom}(E,\mathbb{F}_{2})$ and consider the basis $v_{1},\ldots,v_{n}$ of $E^{\vee}$ dual to the $e_{i}$. For $v\in E^{\vee},$ put $|v|=\#\{i:c_{i}=1\}$ when $v=\sum\limits_{i=1}^{r}c_{i}v_{i}$ with $c_{i}\in\mathbb{F}_{2}$. The representation
$\sigma_{k}=\bigoplus_{|v|=k}v\quad;\quad k=0,1,\ldots,n$
is $S_{n}$-invariant, of degree $\binom{n}{k}$. For a polynomial $f$, write $f]_{i}$ for the coefficient of the degree $i$ term of $f$; in other words so that $f(x)=\sum_{i}[f]_{i}x^{i}$.
Lemma 2.4 ([GJgln], Proposition 2).
For $e\in\mathcal{O}_{k}$, we have
$\chi_{\sigma_{i}}(e)=\left[(1-x)^{k}(1+x)^{n-k}\right]_{i}.$
Moreover, any $S_{n}$-invariant representation $\sigma$ of $E$ is a direct sum of $\sigma_{k}$’s:
(6) $\sigma=\bigoplus_{k=0}^{n}m_{k}(\sigma)\sigma_{k}$
where $m_{k}(\sigma)$ are certain non-negative integers. The coefficients $m_{k}(\sigma)$ can be expressed in terms of character values of $\sigma$ at the elements of $E$.
For $0\leq k\leq n$, put $\vartheta_{k}=e_{1}+\ldots+e_{k}$. (In particular, $\vartheta_{0}=0$.) Then,
(7) $m_{k}(\sigma)=\frac{1}{2^{n}}\sum\limits_{i=0}^{n}\chi_{\sigma_{i}}(\vartheta_{k})\chi_{\sigma}(\vartheta_{i}).$
From Lemma 2.4, the character value $\chi_{\sigma_{i}}(\vartheta_{k})$ is the coefficient of $x^{i}$ in the expression $(1-x)^{k}(1+x)^{n-k}$. Following [GJgln, Theorem 7] we have
(8) $w(\sigma)=\prod\limits_{k=1}^{n}\Big(\prod\limits_{|v|=k}(1+v)\Big)^{m_{k}(\sigma)}.$
For later use, we state the following lemma; its proof is immediate.
Lemma 2.5.
Let $\sigma$ be an $S_{n}$-invariant representation of $E$ such that $\chi_{\sigma}(\vartheta_{i})=\chi_{\sigma}(\vartheta_{n-i})$ for all $0\leq i\leq n$. Then, we have
[…]
for $1\leq k\leq n.$
2.9. Dickson Invariants
With $E$ as the elementary abelian $2$-group above, write $\operatorname{Sym}(E^{\vee})$ for the symmetric algebra of $E^{\vee}$ over $\mathbb{F}_{2}$. In this algebra, we have an element
$\mathcal{D}(E)=\prod_{v\in E^{\vee}}(1+v)=1+\sum_{i=1}^{n}d_{i}(E),$
for certain $\operatorname{GL}(E)$-invariant polynomials $d_{i}$ of degree $2^{n}-2^{n-i}$, known as Dickson invariants. Note that $w(\operatorname{reg}(E))=\mathcal{D}(E)$, where $\operatorname{reg}(E)$ is the regular representation (as mentioned in [wilk]).
Certain factors of $\mathcal{D}(E)$ enter into our work, which we explain here. Consider the basis ${\bm{v}}=\{v_{1},\ldots,v_{n}\}$ of $E^{\vee}$. Identifying $E^{\vee}$ with $\operatorname{H}^{1}(E)$, we may write
$\operatorname{H}^{*}(E)=\operatorname{Sym}(E^{\vee})\cong\mathbb{F}_{2}[v_{1},\ldots,v_{n}].$
Put
$\mathcal{D}^{[k]}({\bm{v}})=\prod_{|v|=k}(1+v)\in\operatorname{Sym}(E^{\vee}),$
so that $\mathcal{D}(E)=\prod_{k}\mathcal{D}^{[k]}({\bm{v}})$. Clearly, $\mathcal{D}^{[k]}({\bm{v}})$ is a symmetric polynomial. We have
$\mathcal{D}^{[n]}({\bm{v}})=1+\mathcal{E}_{1}\quad\quad\text{ and }\quad\quad\mathcal{D}^{[1]}({\bm{v}})=1+\mathcal{E}_{1}+\mathcal{E}_{2}+\ldots+\mathcal{E}_{n},$
For $n=4$, we have
$\begin{split}\mathcal{D}^{[2]}({\bm{v}})&=1+\mathcal{E}_{1}+\mathcal{E}_{1}^{2}+\mathcal{E}_{1}^{3}+(\mathcal{E}_{2}^{2}+\mathcal{E}_{1}\mathcal{E}_{3})+(\mathcal{E}_{1}\mathcal{E}_{2}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{3})+(\mathcal{E}_{1}\mathcal{E}_{2}\mathcal{E}_{3}+\mathcal{E}_{3}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{4})\text{ and }\\ \mathcal{D}^{[3]}({\bm{v}})&=1+\mathcal{E}_{1}+(\mathcal{E}_{1}^{2}+\mathcal{E}_{2})+(\mathcal{E}_{3}+\mathcal{E}_{1}^{3})+(\mathcal{E}_{1}^{2}\mathcal{E}_{2}+\mathcal{E}_{1}\mathcal{E}_{3}+\mathcal{E}_{4}).\end{split}$
Thus we can rewrite (8) as:
Proposition 2.6.
For an $S_{n}$-invariant representation $\sigma$ of $E$, we have
$w(\sigma)=\prod\limits_{k=1}^{n}\mathcal{D}^{[k]}({\bm{v}})^{m_{k}(\sigma)}$
where $m_{k}(\sigma)$ is given by (7).
3. Symplectic Groups
3.1. Subgroups of $G$
Let $\mathcal{S}=\operatorname{SL}(2,q)$. Let $Z=\{\pm{1}\}$ be the center of $\mathcal{S}$. Set
$J=\begin{pmatrix}0&&0&&\ldots&&0&&1\\ 0&&0&&\ldots&&-1&&0\\ \vdots&&\vdots&&\iddots&&\vdots&&\vdots\\ 0&&1&&\ldots&&0&&0\\ -1&&0&&\ldots&&0&&0\\ \end{pmatrix}$,
and put
$G=\operatorname{Sp}(2n,q)=\{A\in\operatorname{GL}(2n,q):A^{t}JA=J\}.$
Let $X$ be the subgroup of matrices in $G$ of the form
$A=\begin{pmatrix}a_{n}&0&0&\ldots&\ldots&0&0&b_{n}\\ 0&\ddots&&&&&\iddots&0\\ &&a_{2}&&&b_{2}&&\\ \vdots&&&a_{1}&b_{1}&&&\vdots\\ \vdots&&&c_{1}&d_{1}&&&\vdots\\ &&c_{2}&&&d_{2}&&\\ 0&\iddots&&&&&\ddots&0\\ c_{n}&0&0&\ldots&\ldots&0&0&d_{n}\\ \end{pmatrix}$,
meaning the nonzero entries of $A$ lie either on the diagonal or the antidiagonal. Let $Z_{X}$ be the center of $X$; it is the subgroup of diagonal matrices in $G$ which have $1$ or 1-1 on the diagonal.
Note that each $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}\in\mathcal{S}$, and $A\mapsto(A_{1},\ldots,A_{n})$ maps $X$ isomorphically to the $n$-fold product $\mathcal{S}^{n}$. It also maps $Z_{X}$ to $Z^{n}$.
Let $M$ be the subgroup of diagonal $n\times n$ block matrices of $G$; then
$\begin{pmatrix}A&0\\ 0&B\end{pmatrix}\mapsto A$
maps $M$ isomorphically to $\operatorname{GL}(n,q)$. The subgroup of $n\times n$ permutation matrices in $\operatorname{GL}(n,q)$ thus gives a copy of $S_{n}$ in $G$; this subgroup normalizes $X$, and acts by conjugation on $Z_{X}<X$ by the given permutation.
3.2. Detecting SWCs
In this Section, we prove Theorem 1.1. Our strategy is to combine the facts, reviewed below, that $X$ is a detecting subgroup of $G$, and that $Z$ detects the SWCs of $\mathcal{S}$. Along the way, we must determine $\operatorname{H}^{*}_{\operatorname{SW}}(\mathcal{S}^{n})$.
It is well known (see [FP, Chapter VI, Sec. 5] for instance) that the mod $2$ cohomology ring of $\mathcal{S}$ is
$\operatorname{H}^{*}(\mathcal{S})\cong\mathbb{F}_{2}[\mathfrak{e}]\otimes\mathbb{F}_{2}[\mathfrak{b}]/\langle\mathfrak{b}^{2}\rangle$
with $\deg(\mathfrak{b})=3$, $\deg(\mathfrak{e})=4$.
In [NSSL2, Corollary 4.8], we specified an orthogonal representation $\eta$ of $\mathcal{S}$ with the property that $w_{4}(\eta)=\mathfrak{e}$. We also studied the restriction map $\operatorname{H}^{*}(\mathcal{S})\to H^{*}(Z)$ and computed that $\mathfrak{e}$ maps to $v^{4}\in\operatorname{H}^{4}(Z)$. This gives an isomorphism
(9) $\operatorname{H}^{4}(\mathcal{S})\overset{\sim}{\to}\operatorname{H}^{4}(Z),$
and leads to:
Theorem 3.1 ([NSSL2]).
The center $Z$ detects SWCs of $\mathcal{S}$. We have $\operatorname{H}^{*}_{\operatorname{SW}}(\mathcal{S})=\mathbb{F}_{2}[\mathfrak{e}]$.
From above, consider the subgroup $X$ of $G=\operatorname{Sp}(2n,q)$, isomorphic to $\mathcal{S}^{n}$. There are projections
$\text{pr}_{j}:\mathcal{S}^{n}\to\mathcal{S}\quad\text{ ; }\quad 1\leq j\leq n,$
and by Künneth we have
(10) $\operatorname{H}^{*}(X)\cong\operatorname{H}^{*}(\mathcal{S}^{n})\cong\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]\otimes_{\mathbb{F}_{2}}\mathbb{F}_{2}[\mathfrak{b}_{1},\ldots,\mathfrak{b}_{n}]/(\mathfrak{b}_{1}^{2},\ldots,\mathfrak{b}_{n}^{2}),$
where $\mathfrak{e}_{j}=\text{pr}^{*}_{j}(\mathfrak{e})$ and $\mathfrak{b}_{j}=\text{pr}^{*}_{j}(\mathfrak{b})$. Note that $\mathfrak{e}_{j}=w_{4}(\eta_{j})$ with $\eta_{j}=\eta\circ\operatorname{pr}_{j}$ for each $j$.
Lemma 3.2 ([Milgram], Chapter VII, Lemma 6.2).
The subgroup $X$ detects the mod $2$ cohomology of $G$.
From before, the subgroup $Z_{X}$ of $G$ is normalized by $S_{n}$ and its mod $2$ cohomology ring is
$\operatorname{H}^{*}(Z_{X})\cong\operatorname{H}^{*}(C_{2}^{n})\cong\mathbb{F}_{2}[v_{1},\ldots,v_{n}]$
by (4).
Consider the case $n=2$. Then by (10) we have $\operatorname{H}^{k}(\mathcal{S}\times\mathcal{S})=0$ when $k\equiv 1\pmod{4}$, and $\operatorname{H}^{k}(\mathcal{S}\times\mathcal{S})\subseteq\mathbb{F}_{2}[\mathfrak{e}_{1},\mathfrak{e}_{2}]$ when $k\equiv 0\pmod{4}$. We have the following lemma:
Lemma 3.3.
The Steenrod square $\operatorname{Sq}^{2}:\operatorname{H}^{k}(\mathcal{S}\times\mathcal{S})\to\operatorname{H}^{k+2}(\mathcal{S}\times\mathcal{S})$ is trivial when $k$ is a multiple of $4$.
Proof.
By the above, it is enough to show that $\operatorname{Sq}^{2}$ vanishes on $\mathbb{F}_{2}[\mathfrak{e}_{1},\mathfrak{e}_{2}]$.
Since $\operatorname{H}^{5}(\mathcal{S})=\operatorname{H}^{6}(\mathcal{S})=\{0\}$, we necessarily have $\operatorname{Sq}^{1}(\mathfrak{e})=\operatorname{Sq}^{2}(\mathfrak{e})=0$. By naturality of $\operatorname{Sq}$ we also have $\operatorname{Sq}^{1}(\mathfrak{e}_{i})=\operatorname{Sq}^{2}(\mathfrak{e}_{i})=0$ for $i=1,2$. By Cartan’s formula (3), we deduce $\operatorname{Sq}^{2}(\mathfrak{e}_{1}^{s}\mathfrak{e}_{2}^{t})=0$ for all nonnegative $s,t$. The conclusion follows. ∎
Proposition 3.4.
We have
$\operatorname{H}^{*}_{\operatorname{SW}}(\mathcal{S}\times\mathcal{S})=\mathbb{F}_{2}[\mathfrak{e}_{1},\mathfrak{e}_{2}].$
Proof.
Since $w_{4}(\eta_{i})=\mathfrak{e}_{i}$ for $i=1,2$, the right side is contained in the left side.
Now let $\pi$ be an orthogonal representation of $\mathcal{S}\times\mathcal{S}$. By (10) again it is enough to see that $w_{k}(\pi)=0$ whenever $k\equiv 2,3$ (mod $4$).
For $k\equiv 2$ (mod $4$), we use Wu’s formula (Proposition 2.1) with $i=2$ and j=k2j=k-2, giving
$\begin{split}\operatorname{Sq}^{2}(w_{k-2}(\pi))&\small={k-5\choose 0}w_{2}(\pi)w_{k-2}(\pi)+{k-4\choose 1}w_{1}(\pi)w_{k-1}(\pi)+{k-3\choose 2}w_{0}(\pi)w_{k}(\pi)\\ &=\frac{(k-3)(k-4)}{2}w_{k}(\pi).\end{split}$
The second equality holds because $\operatorname{H}^{i}(\mathcal{S}\times\mathcal{S})=\{0\}$ for $i=1,2$. Also $\frac{(k-3)(k-4)}{2}$ is odd for $k\equiv 2$ (mod $4$), and $\operatorname{Sq}^{2}(w_{k-2}(\pi))=0$ by Lemma 3.3. Hence $w_{k}(\pi)=0$.
A similar argument works for $k\equiv 3$ (mod $4$): applying Wu’s formula with $i=1$ and j=k1j=k-1 shows that $w_{k}(\pi)=0$. ∎
Theorem 3.5.
We have
$\operatorname{H}^{*}_{\operatorname{SW}}(\mathcal{S}^{n})=\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}].$
Proof.
Again we have $\mathfrak{e}_{i}=w_{4}(\eta_{i})\in\operatorname{H}^{*}_{\operatorname{SW}}(\mathcal{S}^{n})$, so the left hand side contains the right hand side. For the other direction, it is enough to show that the SWCs of an OIR of $\mathcal{S}^{n}$ lie in $\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]$.
First we treat the case of “decomposable” orthogonal representations. For $0\leq r\leq n$ with $r$ even, let $\varphi_{1},\ldots,\varphi_{r}$ be symplectic representations of $\mathcal{S}$, and $\pi_{1},\ldots,\pi_{s}$ orthogonal representations of $\mathcal{S}$, with $r+s=n$. Let
$\Pi=\varphi_{1}\boxtimes\cdots\boxtimes\varphi_{r}\boxtimes\pi_{1}\boxtimes\cdots\boxtimes\pi_{s}$
be their external product. Then $\Pi$ is an orthogonal representation of $\mathcal{S}^{n}$, as in Section 2.6. Note that each of
$\varphi_{1}\boxtimes\varphi_{2}\>,\>\ldots,\>\varphi_{r-1}\boxtimes\varphi_{r}$
is an orthogonal representation of $\mathcal{S}\times\mathcal{S}$. By Theorem 3.1, Proposition 3.4 and Corollary A.1.1, we deduce $w(\Pi)\in\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]$.
Next we treat symmetrizations of decomposable representations. For representations $\pi_{1},\ldots,\pi_{n}$ of $\mathcal{S}$, put $\Pi=\pi_{1}\boxtimes\cdots\boxtimes\pi_{n}$. Then, $w(S(\Pi))=\kappa(c(\Pi))$. From Corollary A.1.1, $c(\Pi)$ lies in the subalgebra of $\operatorname{H}^{*}(\mathcal{S}^{n},\mathbb{Z})$ generated by the $p_{j}^{*}c_{i}(\pi_{j})$. Apply the homomorphism $\kappa$; then Theorem 3.1 gives
$w(S(\Pi))\in\mathbb{F}_{2}[p_{j}^{*}w_{2i}(S(\pi_{j}))]\subseteq\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}].$
By the classification in Section 2.6, up to the permutation action of $S_{n}$ on $\mathcal{S}^{n}$, each OIR takes one of the forms treated above. Since $\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]$ is stable under this action, we are done. ∎
Now we are ready to prove our main detection theorem.
Proof of Theorem 1.1.
From Equation (9), the restriction map $\operatorname{H}_{\operatorname{SW}}^{*}(\mathcal{S}^{n})\to\operatorname{H}^{*}(Z^{n})$ maps $\mathfrak{e}_{i}$ to $v_{i}^{4}$; in our coordinates this gives an isomorphism
$\mathbb{F}_{2}[\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n}]\cong\mathbb{F}_{2}[v_{1}^{4},\ldots,v_{n}^{4}].$
Now by Lemma 3.2, we have
[…]
The image lands in the $S_{n}$-invariant subalgebra, as in Section $2.$ ∎
3.3. Product Formula for Total SWCs
Let $G=\operatorname{Sp}(2n,q)$, and $\pi$ be an orthogonal representation of $G$. To find $w(\pi)$, we work with
$w(\pi|_{Z_{X}})\in\mathbb{F}_{2}[v_{1}^{4},\ldots,v_{n}^{4}]^{S_{n}}$
due to Theorem 1.1. Since $\pi$ is $S_{n}$-invariant, we can write $\pi|_{Z_{X}}\cong\bigoplus\limits_{k=0}^{n}m_{k}(\pi)\sigma_{k}$ (with $\sigma_{k}$ as in Section 2.8) and its total SWC is described by Proposition 2.6. But we can say more about the coefficients $m_{k}(\pi)$ appearing in $w(\pi|_{Z_{X}})$ because $\pi|_{Z_{X}}$ is coming from a representation of the bigger group $G$.
Let $2^{a+1}$ be the highest power of $2$ which divides $|\mathcal{S}|$. Gorenstein in the proof of [goren, Theorem 8.3, Chapter 2] specifies elements $x,y\in\mathcal{S}$ which generate a $2$-Sylow subgroup of $\mathcal{S}$. (It is generalized quaternion of order $2^{a+1}$.) One checks that the subgroup $Q$ generated by $y$ and $x^{2^{a-2}}$ is the quaternion group of order $8$. As $X$ is isomorphic to $\mathcal{S}^{n}$, it correspondingly has a subgroup $Q_{X}$ isomorphic to $Q^{n}$, and containing $Z_{X}$. We thus have a chain of subgroups $Z_{X}\leq Q_{X}\leq X\leq G$. Since $Z_{X}$ detects SWCs of $G$, we infer that $Q_{X}$ also detects the SWCs of $G$.
Let $\pi$ be an orthogonal representation of $G$. Clearly $\operatorname{res}^{G}_{Z_{X}}\pi=\operatorname{res}^{Q_{X}}_{Z_{X}}\operatorname{res}^{G}_{Q_{X}}\pi$ and is $S_{n}$-invariant. Now, Lemma 2.3 gives all $m_{k}(\pi)$ are divisible by $4$.
We can now obtain $w(\pi)$ as its image in $\operatorname{H}^{*}(X)$ by identifying $v_{i}^{4}\in\operatorname{H}^{*}(Z_{X})$ with $\mathfrak{e}_{i}\in\operatorname{H}^{*}(X)$. For $i=0,1,\ldots,n$, let
$g_{i}=\operatorname{diag}(\underbrace{1,\ldots,1}_{n-i},\underbrace{-1,\ldots,-1}_{2i},\underbrace{1,\ldots,1}_{n-i})\in G.$
Theorem 3.6.
The total SWC of an orthogonal representation $\pi$ of $G=\operatorname{Sp}(2n,q)$ is given by
(11) $w(\pi)=\prod_{k=1}^{n}\left(\mathcal{D}^{[k]}(\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n})\right)^{m_{k}(\pi)/4},$
where $m_{k}(\pi)=\cfrac{1}{2^{n}}\sum\limits_{i=0}^{n}\chi_{\sigma_{i}}(g_{k})\chi_{\pi}(g_{i})$.
Note that the character values $\chi_{\sigma_{i}}(g_{k})$ are given by Lemma 2.4.
Example 3.7.
For the regular representation $\operatorname{reg}(G)$, we have
$w(\operatorname{reg}(G))=\mathcal{D}(\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n})^{|G|/2^{n+2}}.$
3.4. Gow’s Formula
In this section, let $G=\operatorname{Sp}(2n,q)$.
Theorem 3.8 ([gow1985real], Theorem 1).
Let $\pi$ be an irreducible self-dual representation of $G$ with central character $\omega_{\pi}$ and Frobenius-Schur Indicator $\varepsilon(\pi)$. Then, we have
(12) $\varepsilon(\pi)=\omega_{\pi}(-\mathbbm{1}).$
We simply call this equality Gow’s formula. For irreducible orthogonal $\pi$ of $G$, it means $\chi_{\pi}(\mathbbm{1})=\chi_{\pi}(-\mathbbm{1})$. This is same as $\chi_{\pi}(g_{0})=\chi_{\pi}(g_{n})$, which leads to
(13) $\chi_{\pi}(g_{i})=\chi_{\pi}(g_{n-i})\quad;\quad 0\leq i\leq n.$
Now, Lemma 2.5 simplifies the exponents $m_{k}(\pi)$ in Theorem 3.6 when $\pi$ is either irreducible orthogonal, or the symmetrization of an irreducible symplectic:
Corollary 3.8.1.
Let $\pi$ be an irreducible orthogonal representation of $G$. Then the total SWC of $\pi$ is
$w(\pi)=\prod_{k=1}^{n}\Big(\mathcal{D}^{[k]}(\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n})\Big)^{m_{k}(\pi)/4}$
where
[…]
Corollary 3.8.2.
Let $\varphi$ be an irreducible symplectic representation of $G$. Then we have
$w(S(\varphi))=\prod_{k=1}^{n}\Big(\mathcal{D}^{[k]}(\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n})\Big)^{m_{k}(\varphi)/2}$
where
[…]
In the formula above, $\left\lceil\cdot\right\rceil$ denotes the ceiling function; in particular
$\left\lceil\frac{n-2}{2}\right\rceil=\begin{cases}\frac{n-2}{2},&\text{if $n$ is even},\\[6.0pt] \frac{n-1}{2},&\text{if $n$ is odd}.\end{cases}$
4. Examples
4.1. The case $n=1$
Let $G=\operatorname{SL}(2,q)$, and $\pi$ an orthogonal representation of $G$. Theorem 3.6 applied for $n=1$ gives $w(\pi)=(1+\mathfrak{e})^{m_{1}(\pi)/4}$, where $m_{1}(\pi)$ in terms of character values at $g_{0}=\mathbbm{1}$, $g_{1}=-\mathbbm{1}$ is:
$m_{1}(\pi)$ $=\frac{1}{2}\big(\chi_{\pi}(\mathbbm{1})-\chi_{\pi}(-\mathbbm{1})\big).$
In particular,
(14) $w_{4}(\pi)=\frac{1}{8}(\deg\pi-\chi_{\pi}(-\mathbbm{1}))\mathfrak{e}.$
When $\pi$ is irreducible orthogonal, $m_{1}(\pi)=0$ by Corollary 3.8.1. Therefore, $w(\pi)$ is trivial for such representations. (Compare [NSSL2, Theorem 1.2].)
4.2. The case $n=2$
Let $g_{0}=\mathbbm{1},g_{1}=\operatorname{diag}(1,-1,-1,1),g_{2}=-\mathbbm{1}\in G.$ With $n=2$ in Theorem 3.6, the total SWC of an orthogonal representation $\pi$ of $G$ is:
(15) $\begin{split}w(\pi)&=\Big(\mathcal{D}^{[1]}(\mathfrak{e}_{1},\mathfrak{e}_{2})\Big)^{m_{1}(\pi)/4}\Big(\mathcal{D}^{[2]}(\mathfrak{e}_{1},\mathfrak{e}_{2})\Big)^{m_{2}(\pi)/4}\\ &=\Big((1+\mathfrak{e}_{1})(1+\mathfrak{e}_{2})\Big)^{m_{1}(\pi)/4}\Big(1+\mathfrak{e}_{1}+\mathfrak{e}_{2}\Big)^{m_{2}(\pi)/4},\end{split}$
where
$m_{1}(\pi)$ $=\frac{1}{4}\big(\chi_{\sigma_{0}}(g_{1})\chi_{\pi}(\mathbbm{1})+\chi_{\sigma_{1}}(g_{1})\chi_{\pi}(g_{1})+\chi_{\sigma_{2}}(g_{1})\chi_{\pi}(-\mathbbm{1})\big)$
$=\frac{1}{4}\big(\chi_{\pi}(\mathbbm{1})-\chi_{\pi}(-\mathbbm{1})\big),\text{ and}$
$m_{2}(\pi)$ $=\frac{1}{4}\big(\chi_{\sigma_{0}}(g_{2})\chi_{\pi}(\mathbbm{1})+\chi_{\sigma_{1}}(g_{2})\chi_{\pi}(g_{1})+\chi_{\sigma_{2}}(g_{2})\chi_{\pi}(-\mathbbm{1})\big)$
$=\frac{1}{4}\big(\chi_{\pi}(\mathbbm{1})-2\chi_{\pi}(g_{1})+\chi_{\pi}(-\mathbbm{1})\big).$
Here, the character values $\chi_{\sigma_{i}}(g_{k})$ are obtained using Lemma 2.4, by expanding the polynomials $(1-y)^{k}(1+y)^{2-k}$ for $k=1,2$. (This proves Theorem 1.3.)
Let $\mathcal{E}_{i}=\mathcal{E}_{i}(\mathfrak{e}_{1},\mathfrak{e}_{2})$. From Equation (15), we deduce
[…]
When $\pi$ is irreducible orthogonal, Corollary 3.8.1 leads to the simplification:
$w(\pi)=\Big(1+\mathfrak{e}_{1}+\mathfrak{e}_{2}\Big)^{m_{2}(\pi)/4}$
where
$m_{2}(\pi)=\frac{1}{2}\big(\chi_{\pi}(\mathbbm{1})-\chi_{\pi}(g_{1})\big).$
To prove Corollary 1.3.1, we need to describe some representations of $G=\operatorname{Sp}(4,q)$ and find some of their character values, which we will do now.
First is a parabolically induced representation given as follows. Let $B$ be the Borel subgroup, consisting of the upper triangular matrices of $G$. Consider the Levi subgroup
$M=\left\{\begin{pmatrix}s&0&0&0\\ 0&\ast&\ast&0\\ 0&\ast&\ast&0\\ 0&0&0&s^{-1}\end{pmatrix}:s\in\mathbb{F}_{q}^{\times}\right\}\cong\mathbb{F}_{q}^{\times}\times\mathcal{S}.$
Let $P$ be the parabolic subgroup generated by $M$ and $B$. We have $|P|=q^{4}(q-1)(q^{2}-1)$. Let $T_{1}$ be the subgroup with elements of the form $\operatorname{diag}(s,1,1,s^{-1})$. Consider a linear character $\chi$ of $\mathbb{F}_{q}^{\times}$ with $\chi(-1)=-1$. Set $\alpha=\chi\circ\operatorname{pr}$, where $\operatorname{pr}$ is the projection map $M\twoheadrightarrow T_{1}$. Then, $\alpha(\operatorname{diag}(-1,1,1,-1))=-1$.
Consider the parabolic induction $\pi_{1}=\operatorname{Ind}_{P}^{G}\alpha$. We have $\deg\pi=(q+1)(q^{2}+1)$, $\chi_{\pi_{1}}(-\mathbbm{1})=-(q+1)(q^{2}+1)$ (as $\alpha$ is odd), and $\chi_{\pi_{1}}(g_{1})=0$. (By the Frobenius formula for induced representations.)
Next is a representation induced from $X$. Let $\psi$ be an odd linear character of $\mathbb{F}_{q^{2}}^{\times}$ with $\psi^{q}\neq\psi$. According to [fulton, Section 5.2], there is an irreducible cuspidal representation $\sigma_{\psi}$ of $\mathcal{S}$ corresponding to $\psi$, with degree q1q-1. Viewing $X$ as the product $\mathcal{S}\times\mathcal{S}$ as in Section 3.1, we form the external tensor product representation $\sigma_{\psi}\boxtimes 1$ of $X$. Finally, let $\pi_{2}=\operatorname{Ind}_{X}^{G}(\sigma_{\psi}\boxtimes 1)$. Here, $\deg(\pi_{2})=q^{2}(q-1)(q^{2}+1)$ and again as $\psi$ is odd, $\chi_{\pi_{2}}(-\mathbbm{1})=-\deg\pi_{2}.$ Also, $\chi_{\pi_{2}}(g_{1})=0$: observe that the only conjugates of $g_{1}$ in $X$ are $\pm g_{1}$ and there is a Weyl element sending $g_{1}$ to $g_{1}$ upon conjugation. (And again use the Frobenius character formula.)
Proof of Corollary 1.3.1.
Let $\Pi_{1}=S(\pi_{1})$, $\Pi_{2}=S(\pi_{2})$ with $\pi_{1}$, $\pi_{2}$ as above. Using their character values, we obtain $m_{1}(\Pi_{1})=(q+1)(q^{2}+1)$, $m_{1}(\Pi_{2})=q^{2}(q-1)(q^{2}+1)$ and $m_{2}(\Pi_{2})=m_{2}(\Pi_{2})=0$. Note that $m_{1}(\Pi_{1})/4$ is odd when $q\equiv 1$ mod ($4$), and $m_{1}(\Pi_{2})/4$ is odd when $q\equiv 3$ mod ($4$).
Corresponding to the cases, we have
$w(\Pi_{i})$ $=(1+\mathfrak{e}_{1}+\mathfrak{e}_{2}+\mathfrak{e}_{1}\mathfrak{e}_{2})^{m_{1}(\Pi_{i})/4}$
$=1+\frac{m_{1}(\Pi_{i})}{4}\left(\mathfrak{e}_{1}+\mathfrak{e}_{2}+\mathfrak{e}_{1}\mathfrak{e}_{2}\right)+\ldots$
hence $\mathcal{E}_{1},\mathcal{E}_{2}\in\operatorname{H}^{*}_{\operatorname{SW}}(G)$.
4.3. The case $n=4$
Let $\underline{\mathfrak{e}}=(\mathfrak{e}_{1},\mathfrak{e}_{2},\mathfrak{e}_{3},\mathfrak{e}_{4})$. We apply Theorem 3.6 for $n=4$ to have the total SWC of an orthogonal representation $\pi$ of $G$:
(18) $w(\pi)=\mathcal{D}^{[1]}(\underline{\mathfrak{e}})^{m_{1}(\pi)/4}\mathcal{D}^{[2]}(\underline{\mathfrak{e}})^{m_{2}(\pi)/4}\mathcal{D}^{[3]}(\underline{\mathfrak{e}})^{m_{3}(\pi)/4}\mathcal{D}^{[4]}(\underline{\mathfrak{e}})^{m_{4}(\pi)/4};$
please see Section 2.9 for calculation of these $\mathcal{D}^{[i]}(\underline{\mathfrak{e}})$.
Also, Gow’s formula through Corollary 3.8.1 allows simplification for irreducible orthogonal $\pi$:
$w(\pi)=\Big(1+\mathcal{E}_{1}+\mathcal{E}_{1}^{2}+\mathcal{E}_{1}^{3}+\mathcal{E}_{2}^{2}+\mathcal{E}_{1}\mathcal{E}_{3}+\mathcal{E}_{1}\mathcal{E}_{2}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{3}+\mathcal{E}_{1}\mathcal{E}_{2}\mathcal{E}_{3}+\mathcal{E}_{3}^{2}+\mathcal{E}_{1}^{2}\mathcal{E}_{4}\Big)^{m_{2}(\pi)/4}(1+\mathcal{E}_{1})^{m_{4}(\pi)/4}$
where, with the help of Lemma 2.4 once again, gives
$\begin{split}m_{2}(\pi)&=\cfrac{1}{8}\big(\deg(\pi)-\chi_{\pi}(g_{2})\big),\\ m_{4}(\pi)&=\cfrac{1}{8}\big(\deg(\pi)-4\chi_{\pi}(g_{1})+3\chi_{\pi}(g_{2})\big).\end{split}$
Again, $g_{1}=\operatorname{diag}(1,1,1,-1,-1,1,1,1)$ and $g_{2}=\operatorname{diag}(1,1,-1,-1,-1,-1,1,1)$.
4.4. Weil Representations
For a fixed nontrivial linear character $\lambda$ of $(\mathbb{F}_{q},+)$, the prescription
$x\mapsto\lambda_{x}\quad;\quad\lambda_{x}(y)=\lambda(xy)$
defines a group isomorphism $\mathbb{F}_{q}\overset{\sim}{\to}\widehat{\mathbb{F}}_{q}$. Let $\alpha$ be a non-square, and $1$ be the identity in $\mathbb{F}_{q}^{\times}$. According to Gérardin [PGWeil, Theorem 2.4(d)], the symplectic group $G=\operatorname{Sp}(2n,q)$ has two nonisomorphic Weil representations $\mathcal{W}_{\lambda_{1}}$, $\mathcal{W}_{\lambda_{\alpha}}$ associated to the characters $\lambda_{1}$, $\lambda_{\alpha}$. Set $\mathcal{W}=\mathcal{W}_{\lambda_{1}}$ and $\mathcal{W}^{\prime}=\mathcal{W}_{\lambda_{\alpha}}$. Both $\mathcal{W}$, $\mathcal{W}^{\prime}$ are complex representations of degree $q^{n}$, and decompose into two irreducible representations as
$\mathcal{W}=\mathcal{W}_{l}\oplus\mathcal{W}_{s}\quad,\quad\mathcal{W}^{\prime}=\mathcal{W}^{\prime}_{l}\oplus\mathcal{W}^{\prime}_{s}$
where $\mathcal{W}_{l}$, $\mathcal{W}^{\prime}_{l}$ are the components with larger degrees $(q^{n}+1)/2$, and $\mathcal{W}_{s}$, $\mathcal{W}^{\prime}_{s}$ have degrees $(q^{n}-1)/2$. When $q\equiv 3$ (mod $4$), none of these four irreducible representations are self-dual. In fact, $\mathcal{W}_{l}$, $\mathcal{W}^{\prime}_{l}$ are dual to each other, and similarly $\mathcal{W}_{s}$ is the dual of $\mathcal{W}^{\prime}_{s}$. Whereas when $q\equiv 1$ (mod $4$), all representations of $G$ are self-dual. In this case, exactly one of $\mathcal{W}_{l}$, $\mathcal{W}_{s}$ is orthogonal, and the other is symplectic. The same is true for the components of $\mathcal{W}^{\prime}$. So, neither $\mathcal{W}$ nor $\mathcal{W}^{\prime}$ is orthogonal. Here we compute
$w(S(\mathcal{W}))=c(\mathcal{W})\mod 2.$
From[PGWeil, Corollary 4.8.1], the character values of $\mathcal{W}$ at $g_{i}$ are given by:
$\chi_{\mathcal{W}}(g_{i})=(-1)^{i\frac{q-1}{2}}q^{N(g_{i})},\text{ with }\quad N(g)=\frac{1}{2}\dim_{\mathbb{F}_{q}}\ker(g-\mathbbm{1}).$
One sees that
$\chi_{\mathcal{W}}(g_{i})=(-1)^{i\frac{q-1}{2}}q^{n-i}.$
For each $1\leq k\leq n$, Equation (7) gives
$m_{k}(S(\mathcal{W}))$ $=\frac{1}{2^{n-1}}\sum_{i=0}^{n}[f_{k}]_{i}\chi_{\mathcal{W}}(g_{i})$
$=\frac{1}{2^{n-1}}\sum_{i=0}^{n}[f_{k}]_{i}(-1)^{i\frac{q-1}{2}}q^{n-i}$
where $f_{k}]_{i}$ is the coefficient of $x^{i}$ in $f_{k}(x)=(1-x)^{k}(1+x)^{n-k}$. Note that $f_{k}]_{n-i}=(-1)^{k}[f_{k}]_{i}$ for each $i=0,\ldots,n$, so that
$m_{k}(S(\mathcal{W}))=\frac{1}{2^{n-1}}\sum_{i=0}^{n}(-1)^{k}[f_{k}]_{n-i}(-1)^{i\frac{q-1}{2}}q^{n-i}.$
Put $F_{k}(x):=(x-1)^{k}(x+1)^{n-k}=(-1)^{k}f_{k}(x)$. Then, for $q\equiv 1$ (mod $4$), it is clear that
$m_{k}(S(\mathcal{W}))=\frac{1}{2^{n-1}}F_{k}(q)=\frac{1}{2^{n-1}}(q-1)^{k}(q+1)^{n-k}.$
Whereas for $q\equiv 3$ (mod $4$), with a few manipulations, we get
$m_{k}(S(\mathcal{W}))$ $=\frac{1}{2^{n-1}}\sum_{i=0}^{n}(-1)^{i}[F_{k}]_{n-i}q^{n-i}$
$=\frac{1}{2^{n-1}}\sum_{j=0}^{n}(-1)^{n-j}[F_{k}]_{j}q^{j}$
$=\frac{(-1)^{n}}{2^{n-1}}F_{k}(-q)$
$=\frac{1}{2^{n-1}}(q-1)^{n-k}(q+1)^{k}.$
Corollary 4.0.1.
For the Weil representations $\mathcal{W}$ and $\mathcal{W}^{\prime}$ of $G$, we have
$w(S(\mathcal{W}))=w(S(\mathcal{W}^{\prime}))=\prod_{k=1}^{n}\mathcal{D}^{[k]}(\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n})^{m_{k}(S(\mathcal{W}))/4}$
where
$m_{k}(S(\mathcal{W}))=\begin{cases}\cfrac{1}{2^{n-1}}(q-1)^{k}(q+1)^{n-k},&q\equiv 1\;(\operatorname{mod}4)\\ \cfrac{1}{2^{n-1}}(q-1)^{n-k}(q+1)^{k},&q\equiv 3\;(\operatorname{mod}4).\end{cases}$
Proof.
We have already shown this for $\mathcal{W}$. In [PGWeil], we see that $\chi_{\mathcal{W}}(g_{i})=\chi_{\mathcal{W}^{\prime}}(g_{i})$ for each $i$. Thus, the formula also holds for $w(S(\mathcal{W}^{\prime}))$. ∎
5. Universal SWCs
Let us write $X_{n}$ for the subgroup of $G_{n}=\operatorname{Sp}(2n,q)$ isomorphic to $\mathcal{S}^{n}$ (previously just written as $X$). The restriction map $\operatorname{H}^{*}(G_{n})$ to $\operatorname{H}^{*}(X_{n})$ has image in the $S_{n}$-invariants $\operatorname{H}^{*}(X_{n})^{S_{n}}$, which contains the elements $\mathcal{E}_{j}=\mathcal{E}_{j}(\mathfrak{e}_{1},\ldots,\mathfrak{e}_{n})$ but also
$\mathcal{F}_{j}=\sum_{\begin{subarray}{c}i_{1}<\ldots<i_{j}\\ 1\leq k\leq j\end{subarray}}\mathfrak{e}_{i_{1}}\ldots\hat{\mathfrak{e}}_{i_{k}}\ldots\mathfrak{e}_{i_{j}}b_{i_{k}}.$
Note that $\deg(\mathcal{E}_{j})=4j$ and $\deg(\mathcal{F}_{j})=4j-1$.
Theorem 5.1.
Let $E_{1},\ldots,E_{n},F_{1},\ldots,F_{n}$ be formal indeterminates. The homomorphism
$\mathbb{F}_{2}[E_{1},\ldots,E_{n},F_{1},\ldots,F_{n}]\to\operatorname{H}^{*}(X_{n})^{S_{n}}$
defined by sending $E_{i}$ to $\mathcal{E}_{i}$ and $F_{i}$ to $\mathcal{F}_{j}$ is surjective with kernel equal to the ideal $(F_{1}^{2},\ldots,F_{n}^{2})$.
Proof.
This is implicit in [FP, Theorem 6.1, page 283]. ∎
Let $m\geq n$ and $\iota_{n}:G_{n-1}\to G_{m}$ be the following inclusion: For $A\in G_{n-1}$,
$\iota_{n}(A)=\begin{pmatrix}\mathbbm{1}_{m-n+1}&&\\ &A&\\ &&\mathbbm{1}_{m-n+1}\end{pmatrix}$.
Proof of Theorem 1.4.
By Lemma 3.2, it is enough to show that the restriction
$\operatorname{H}^{i}(X_{m})^{S_{m}}\to\operatorname{H}^{i}(X_{n-1})^{S_{n-1}}$ is injective. Write $R_{n}=P_{n}\otimes_{\mathbb{F}_{2}}Q_{n}$, where $P_{n}=\mathbb{F}_{2}[\mathcal{E}_{1},\ldots,\mathcal{E}_{n}]$ is the subalgebra generated by $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$ and $Q_{n}=\mathbb{F}_{2}[\mathcal{F}_{1},\ldots,\mathcal{F}_{n}]$ is the subalgebra generated by $\mathcal{F}_{1},\ldots,\mathcal{F}_{n}$ in $\operatorname{H}^{*}(\mathcal{S}^{n})$. From above, we need to show that the restriction of the map $P_{m}\otimes Q_{m}\to P_{n-1}\otimes Q_{n-1}$ to the degree $k$ part is injective for all k<4n1k<4n-1.
Any monomial of degree $k$ is of the form $\mathcal{E}_{1}^{r_{1}}\ldots\mathcal{E}_{m}^{r_{m}}\mathcal{F}_{1}^{s_{1}}\ldots\mathcal{F}_{m}^{s_{m}}$ with $r_{i}\geq 0$, $s_{i}\in\{0,1\}$. We may write this as $\mathcal{E}^{\mathbf{r}}\mathcal{F}^{\mathbf{s}}$, with $\mathbf{r}=(r_{1},\ldots,r_{m})$ and $\mathbf{s}=(s_{1},\ldots,s_{m})$. By Theorem 5.1 above, these monomials are linearly independent. The condition
$\sum_{i=1}^{m}(4i)r_{i}+\sum_{j=1}^{m}(4j-1)s_{j}=k<4n-1$
ensures that if $i\geq n$, then $r_{i}=0$ and $s_{i}=0$.
An element $\alpha=\sum_{(\mathbf{r},\mathbf{s})}c_{\mathbf{r},\mathbf{s}}\mathcal{E}^{\mathbf{r}}\mathcal{F}^{\mathbf{s}}\in\operatorname{H}^{k}(G_{m})$ under $\iota^{*}_{n}$ gets mapped to
$\sum_{\begin{subarray}{c}\mathbf{r}=(r_{1},\ldots,r_{n-1})\\ \mathbf{s}=(s_{1},\ldots,s_{n-1})\end{subarray}}c_{\mathbf{r},\mathbf{s}}\mathcal{E}^{\mathbf{r}}\mathcal{F}^{\mathbf{s}}\in\operatorname{H}^{k}(G_{n-1}).$
If $i^{*}_{n}(\alpha)=0$, then all $c_{\mathbf{r},\mathbf{s}}$ are zero since $\mathcal{E}^{\mathbf{r}}\mathcal{F}^{\mathbf{s}}$ are linearly independent. This implies $\alpha=0$, and hence we have the injectivity of $\iota^{*}_{n}$ for k<4n1k<4n-1. ∎
Proof of Theorem 1.2.
By Theorem 1.4, the map $i^{*}_{2}:\operatorname{H}^{4}(G_{m})\to\operatorname{H}^{4}(G_{1})=\operatorname{H}^{4}(\mathcal{S})$ is injective for all $m\geq 1$. So the formula for $w_{4}(\pi)$ follows from (14). Similarly, since $i^{*}_{3}:\operatorname{H}^{8}(G_{m})\to\operatorname{H}^{8}(G_{2})$ is injective for all $m\geq 3$, we have $w_{8}(\pi)$ from Equation (17). ∎
Appendix A
In this Appendix we discuss the formula for CCs and SWCs of external tensor products of representations. This formula involves an interesting family of polynomials which we now define.
A.1. The Polynomials $\operatorname{P}_{m,n}$
Given non-negative integers $m,n$, we define $q_{m,n}$ be the following polynomial in $\Lambda=\mathbb{Z}[x_{1},\ldots,x_{m},y_{1},\ldots,y_{n}]$:
(19) $q_{m,n}(\mathbf{x},\mathbf{y})=\prod_{i=1}^{m}\prod_{j=1}^{n}(1+x_{i}+y_{j}).$
It is symmetric in the $x_{i}$ and the $y_{j}$ separately, so by the fundamental theorem of symmetric polynomials [b47, Chapter 4, Section 6, Theorem 1], there is a unique polynomial $\operatorname{P}_{m,n}\in\Lambda$ such that
(20) $q_{m,n}(\mathbf{x},\mathbf{y})=\operatorname{P}_{m,n}(\mathcal{E}_{1}(\mathbf{x}),\ldots,\mathcal{E}_{m}(\mathbf{x}),\mathcal{E}_{1}(\mathbf{y}),\ldots,\mathcal{E}_{n}(\mathbf{y})).$
For example, we have $\operatorname{P}_{m,0}({\bf x},{\bf y})=1+\sum_{i=1}^{m}x_{i}$.
A.2. Characteristic Classes of External Tensor Products
Let $G_{1},G_{2}$ be finite groups, and let $p_{i}:G_{1}\times G_{2}\to G_{i}$ be the projection maps. Given complex $G_{i}$-representations $(\pi_{i},V_{i})$, we consider their external tensor product $(\pi_{1}\boxtimes\pi_{2},V_{1}\otimes V_{2})$.
Proposition A.1.
Suppose $(\pi_{1},V_{1}),$ $(\pi_{2},V_{2})$ have respective degrees $m,n$. Then,
$c(\pi_{1}\boxtimes\pi_{2})=\operatorname{P}_{m,n}(p_{1}^{*}c_{1}(\pi_{1}),\ldots,p_{1}^{*}c_{m}(\pi_{1}),p_{2}^{*}c_{1}(\pi_{2}),\ldots,p_{2}^{*}c_{n}(\pi_{2})),$
where $\operatorname{P}_{m,n}$ is as defined above. When $\pi_{1},\pi_{2}$ are orthogonal, then
$w(\pi_{1}\boxtimes\pi_{2})=\operatorname{P}_{m,n}(p_{1}^{*}w_{1}(\pi_{1}),\ldots,p_{1}^{*}w_{m}(\pi_{1}),p_{2}^{*}w_{1}(\pi_{2}),\ldots,p_{2}^{*}w_{n}(\pi_{2})).$
We will sketch the proof for SWCs in this Appendix; the case of CCs is similar. Iterating Proposition A.1 gives:
Corollary A.1.1.
Let $\Pi=\pi_{1}\boxtimes\cdots\boxtimes\pi_{n}$ be a (complex) representation of $G=G_{1}\times\cdots\times G_{n}$. Then $c_{k}(\Pi)$ is in the subalgebra of $\operatorname{H}^{*}(G,\mathbb{Z})$ generated by $\{p_{j}^{*}c_{i}(\pi_{j}):1\leq j\leq n,i\geq 0\}$. When each $\pi_{i}$ is orthogonal, the class $w_{k}(\Pi)$ is in the subalgebra of $\operatorname{H}^{*}(G)$ generated by $\{p_{j}^{*}w_{i}(\pi_{j}):1\leq j\leq n,i\geq 0\}$.
A.3. Products of Vector Bundles
The results in the previous section are valid for vector bundles over paracompact spaces. Here, we will prove them in this more general setting. All our base spaces are paracompact.
Let $E_{1}$, $E_{2}$ be two real vector bundles over a base space $B$. Their internal tensor product $E_{1}\otimes E_{2}$ is again a real vector bundle over $B$.
Proposition A.2 ([milnor], Chapter 7, Problem 7-C).
Let $E_{1}$, $E_{2}$ be real vector bundles over $B$ with respective ranks $m,n$. Then,
$w(E_{1}\otimes E_{2})=\operatorname{P}_{m,n}(w_{1}(E_{1}),\ldots,w_{m}(E_{1}),w_{1}(E_{2}),\ldots,w_{n}(E_{2})).$
Let $E_{1}$, $E_{2}$ be real vector bundles over base spaces $B_{1}$, $B_{2}$ respectively with $\Pi_{i}:E_{i}\to B_{i}$. We can form their external tensor product $E_{1}\boxtimes E_{2}$ over $B_{1}\times B_{2}$ as follows.
Consider the projection maps $p_{i}:B_{1}\times B_{2}\to B_{i}$. Let $p_{i}^{*}E_{i}$ be the pullback of $E_{i}$ by $p_{i}$ consisting of elements $((b_{1},b_{2}),e_{i})\in(B_{1}\times B_{2})\times E_{i}$ such that $\Pi_{i}(e_{i})=p_{i}(b_{1},b_{2})=b_{i}$. These are vector bundles over the same base space $B_{1}\times B_{2}$. Thus, we construct their internal tensor product, and put $E_{1}\boxtimes E_{2}:=p_{1}^{*}E_{1}\otimes p_{2}^{*}E_{2}$, which is again a vector bundle over $B_{1}\times B_{2}.$ This with the naturality of SWCs and Proposition A.2 gives:
Proposition A.3.
Let $E_{1}$, $E_{2}$ be real vector bundles over respective base spaces $B_{1}$, $B_{2}$ with $\operatorname{rank}(E_{1})=m$ and $\operatorname{rank}(E_{2})=n$. Then,
$w(E_{1}\boxtimes E_{2})=\operatorname{P}_{m,n}(p_{1}^{*}w_{1}(E_{1}),\ldots,p_{1}^{*}w_{m}(E_{1}),p_{2}^{*}w_{1}(E_{2}),\ldots,p_{2}^{*}w_{n}(E_{2})).$
A.4. Vector Bundles From Representations
Let $G$ be a finite group, and $(\rho,U)$ be a real representation of $G$. Associated to $G$ is a classifying space $BG$ with a contractible right principal $G$-bundle EGEG. From $(\rho,U)$ one can form the associated real vector bundle $EG[U]$ over $BG$. To define $EG[U]$, first form the product $EG\times U$; it is $G$-space under the action $g\cdot(x,u)=(x\cdot g^{-1},\rho(g)u)$. Then $EG[U]$ is the quotient of $EG\times U$ by this action. Then put $w_{i}^{\mathbb{R}}(\rho)=w_{i}(EG[U])$. (See for instance [Benson] or [GKT].) The singular cohomology $\operatorname{H}^{*}(BG,A)$ is naturally isomorphic to the group cohomology $\operatorname{H}^{*}(G,A)$ for any abelian group $A$.
Let $(\pi,V)$ be a complex orthogonal representation of $G$. There is a representation $(\rho,U)$, with $U$ a real vector space, so that $\rho\otimes_{\mathbb{R}}\mathbb{C}\cong\pi$ (see [BrokerDieck, Chapter II, Section 6] for instance). Such a representation is called a real form of $\pi$; it is unique up to isomorphism. We now can define $w(\pi):=w^{\mathbb{R}}(\rho)$.
Let $G_{1},G_{2}$ be finite groups with orthogonal complex representations $(\pi_{1},V_{1})$ and $(\pi_{2},V_{2})$. Let $G=G_{1}\times G_{2}$. We can form their external tensor product $(\pi_{1}\boxtimes\pi_{2},V_{1}\otimes V_{2})$, an orthogonal representation of $G$. Let $(\rho_{i},U_{i})$ be real forms of $\pi_{i}$ for $i=1,2$. Then $\rho_{1}\boxtimes\rho_{2}$ is a real form of $\pi_{1}\boxtimes\pi_{2}$. From above, one has
[…]
where $\mathcal{U}_{i}=EG_{i}[U_{i}]$. We may identify $BG_{1}\times BG_{2}$ with $BG$. Now, Proposition A.1 for SWCs follows from Proposition A.3 with $E_{i}=\mathcal{U}_{i}$.
Acknowledgments.
Part of this paper comes out of the first author’s Ph.D. thesis [Malik.thesis] at IISER Pune, during which she was supported by a Ph.D. fellowship from the Council of Scientific and Industrial Research, India. Both authors thank the Chennai Mathematical Institute for its hospitality during many visits.